OpenStudy (dehelloo):

Matrix A is a 1 x 3 Matrix B is a 3 x 1 Matrix Which one of the follow matrix expressions involving A and B is defined? 2B x 3A or A^2 B ........................... It seems to me both are feasible But apparently A^2B is wrong according to the textbook

OpenStudy (anonymous):

mxn matrix can multiply an nxr matrix

OpenStudy (dehelloo):

No idea what that means.

OpenStudy (anonymous):

its A^2 B

OpenStudy (anonymous):

oooops did it wrong sry

OpenStudy (anonymous):

nevermind that

OpenStudy (anonymous):

for example 3x2 * 2x7 matrix is defined, because the 2's match

OpenStudy (anonymous):

a mxn matrix has m rows, n columns a n x r matrix has n rows and r columns

OpenStudy (dehelloo):

yeah so if a= 4x4 a^2= 1x3 * 1x3 So (A) 1x3 x (B) 3x1 also match So cant both be expressed?

OpenStudy (dehelloo):

disregard the 4x4 sorry

OpenStudy (dehelloo):

so both can be expressed but apparently the answer is only 2B x 3A

OpenStudy (anonymous):

lets make a notation [3x1] means a matrix that has dimensions 3 rows and 1 column

OpenStudy (anonymous):

I will use * to indicate matrix multiplication 2B * 3A : 2[3x1] * 3[1x3] = [3x1 ] * [1x3] multiplying a matrix by a scalar does not change dimensions this is defined

OpenStudy (anonymous):

A^2 * B : [1x3 ] * [1x3] * [ 3 x 1] here we see it is not defined because of the 3 and 1 A*A is not defined

OpenStudy (anonymous):

A^2 B = A*A*B but A*A is not defined , so the entire product cannot be defined

OpenStudy (dehelloo):

So the product of 2B * 3A = 3x3 But A^2 = 1x1 So cant both be defined?

OpenStudy (anonymous):

no , A^2 is not 1x1, its not defined at all

OpenStudy (anonymous):

dimension mismatch error

OpenStudy (dehelloo):

I meant A^2 x B = 1x1

OpenStudy (dehelloo):

Sorry

OpenStudy (anonymous):

you cant multiply A by A

OpenStudy (anonymous):

nope, A^2 * B is not defined. you cant multiply an undefined matrix by something else

OpenStudy (anonymous):

A^2 being undefined

OpenStudy (dehelloo):

yeah thanks understand now

OpenStudy (anonymous):

you should do the multiplication from left to right A^2B = AAB